OpenStudy (anonymous):

MEDAL AND FANS :) A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

-4x+16y=12 would this equation count for part a?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

how did you get your new equation?

OpenStudy (anonymous):

i got -4 because it is a multiple of -16 and 16 is the sum of one of the equations i just plugged them in

OpenStudy (anonymous):

@phi are you still there?

OpenStudy (phi):

call the first equation A and the second equation B they want you to replacing one equation with the sum of that equation and a multiple of the other. in other words, if you pick A to replace, they want you to find a new equation (call it C) equal to A + nB where n is a number. if you pick n=1 that makes it easy: the new equation C= A+B can you do that ?

OpenStudy (anonymous):

yeah thanks is that all for part a?

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

umm hold on

OpenStudy (anonymous):

-4+20=16

OpenStudy (phi):

the equations have variables x and y they (usually) never go away. here is how to add the two equations: -3x + 7y = -16 -9x + 5y = 16 you add the left hand sides and the right hand sides: -3x + 7y + -9x + 5y = -16+16

OpenStudy (phi):

you can simplify that. For example notice you have 7 y's and 5 more y's how many y's do you have on the left hand side?

OpenStudy (anonymous):

12

OpenStudy (phi):

you write 12y (remember we are counting how many y's) we also have -3x - 9x can you simplify that ?

OpenStudy (anonymous):

-12x

OpenStudy (phi):

so you now have -12x + 12y = -16+16 simplify the right side: -16+16 is 0. we get -12x + 12y = 0 notice if we start with -3x + 7y = -16 -9x + 5y = 16 we could just "add the columns" -12x + 12y = 0

OpenStudy (anonymous):

thank you so much :) can you also help with part b

OpenStudy (phi):

part B uses the same idea of adding equations. we have to solve the problem two ways. The first way is with the original equations -3x + 7y = -16 -9x + 5y = 16 try this: multiply the first equation by -3. that means multiply both sides and *all terms* each by -3. can you do that ?

OpenStudy (anonymous):

yeah 9x+-21y=-48

OpenStudy (phi):

-3*-16 is not -48

OpenStudy (anonymous):

oh positive sorry

OpenStudy (phi):

yes. when we do these there are painful if you make a simple mistake. they take a lot of work

OpenStudy (anonymous):

yes!!!!!!

OpenStudy (phi):

so we now have these equations (the first multiplied by -3, the 2nd as it was) 9x + -21y= 48 -9x + 5y = 16 add the two equations. what do you get ?

OpenStudy (anonymous):

-16y=64

OpenStudy (anonymous):

is that all @phi

OpenStudy (phi):

notice that the 9x + -9x added to zero, and the x "went away" (that is why I picked -3 to multiply the first equation ... I wanted a +9x) to solve -16y=64 divide both sides by -16 what do you get?

OpenStudy (anonymous):

4 yay thank you

OpenStudy (phi):

you seem to have trouble with the minus signs. make a note: be careful of the signs

OpenStudy (phi):

-16x/-16 = 64/-16 -16/-16 is 1 x = 64/-16 x= -4

OpenStudy (anonymous):

leave it up to me to make a small mistake but thank you so so much you are a lifesavor

OpenStudy (phi):

except that should be y= -4 ! we use that in either of the equations to find x for example -3x + 7y = -16 replace y with -4 -3x + 7*-4 = -16 can you solve for x ?

OpenStudy (anonymous):

-6.3.....

OpenStudy (anonymous):

is that right @phi

OpenStudy (anonymous):

or is it -4?? @phi

OpenStudy (phi):

-3x + 7*-4 = -16 7 * -4 = -28 -3x -28 = -16 add +28 to both sides -3x -28 + 28 = -16 + 28 -3x = 12 divide both sides by -3 x = 12/-3 = -4 yes, x= -4, so the solution is x=-4, y=-4

OpenStudy (anonymous):

thank you so much *hugs you with death grip* sorry for bugging you you saved my life

OpenStudy (phi):

Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points) the new system (as you may remember) is -12x + 12y = 0 -9x + 5y = 16 one way to show this system has the same solution, is to use x=-4 and y=-4 and show that *both* equations are true. can you do that ?

OpenStudy (anonymous):

yeah thats all you help me so much thank you again